# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v4

@@ 1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our proposed method (0.472 vs. 0.809 WinMass). This sensitivity arises because DPO’s objective only regularizes the difference in log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
To address these limitations, we introduce Anchored Direct Preference Optimization (ADPO), a unified framework that reveals a fundamental connection across seemingly disparate learning paradigms. Through a simple anchored projection formulation—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ where $\tilde{p}_{\theta}$ is defined via anchored logits $(s-s^{\mathrm{ref}})/\tau$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all emerge as special cases through different choices of target distribution $q$, anchor policy $\pi_{\mathrm{ref}}$, and temperature $\tau$ (Section 3, Proposition 3.1). The anchoring structure is key: by centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
We prove that this anchoring mechanism provides a principled geometric framework: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.3), yielding an implicit trust region centered at the target (not the anchor itself). This provides a robust, distribution-space analogue to trust-region methods like TRPO [23], without requiring explicit constraints. Crucially, this geometric foundation holds regardless of anchor choice, enabling us to investigate the role of the anchor update strategy. To our knowledge, this is the first systematic comparison of a fixed anchor (frozen throughout training) versus a dynamic anchor (periodically updated to track the current policy) in the context of preference learning and policy optimization. Our central finding is a crucial task-dependent trade-off:
  • A dynamic anchor excels at online exploration in noisy environments, where its adaptive trust region improves performance by 5–11% in 3 out of 4 scenarios.
@@ 2 Related Work
Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the Euclidean probability space. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates. Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in a coordinate system whose origin is fixed at the reference. This anchoring keeps the Fisher metric centered and effectively forms a “quadratic bowl” around the reference distribution, yielding an implicit trust region without any explicit KL constraint. To our knowledge, no prior work has formulated such an anchored cross-entropy that maintains a fixed geometric center in probability space.
@@
$\max_{\pi}\ \mathbb{E}_{x\sim\mathcal{D}}\Big[\mathbb{E}_{\tau\sim\pi(\cdot|x)}r(x,\tau)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|x)\big)\Big],$
@@
$\pi^{*}(\tau|x)\;=\;\frac{1}{Z(x)}\exp\!\Big(\tfrac{1}{\alpha}\,r(x,\tau)\Big),$ (1)
@@
$r(x,\tau)\;=\;\alpha\,\log\pi^{*}(\tau|x)\;+\;\alpha\,\log Z(x),$ (2)
so any relative quantity (pairwise/listwise) cancels $Z(x)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs. (3)–(6), (20)–(27)).
@@
If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [18], Prop. 3.1 and App. D.2.
@@ 3 Anchored Direct Preference Optimization
Throughout this section, we use the following notation:
  • $x$: context (prompt/state); $S_{x}=\{1,\ldots,|S_{x}|\}$: candidate set (actions/outputs)
  • $\pi_{\theta}(i\mid x)$: student policy probability for candidate $i$ given context $x$
  • $\pi_{\mathrm{ref}}(i\mid x)$: reference (anchor) policy, either fixed or dynamic
  • $s_{i}=\log\pi_{\theta}(i\mid x)$, $s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x)$: log-probabilities
  • $q(i\mid S_{x})$: target preference distribution over candidates (from teacher, rewards, or labels)
  • $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$: anchored logits
  • $\tilde{p}_{\theta}(i\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution
@@
$\tau$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau$ (listwise)
@@
Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them. We set $\beta=\beta_{r}=1.0$ by default. However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau}(\tilde{p}-q)$). Thus smaller $\beta$ slows pairwise updates; smaller $\tau$ tightens the trust region via curvature $\propto 1/\tau^{2}$ but increases gradient magnitudes.
@@
Remark (temperature $\tau$ and regularization strength): In listwise settings with anchoring, $\tau$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ (Proposition 3.2). Intuitively:
@@
  • As $\tau\to 0$: $q^{\tau}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
@@
  • As $\tau\to\infty$: $q^{\tau}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
@@
Note that $\tilde{p}_{\theta}$ always matches $q$ at optimality regardless of $\tau$; the temperature affects how the underlying policy $\pi_{\theta}$ relates to the anchor. Smaller $\tau$ induces a tighter implicit trust region in the sense that large deviations $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are more heavily penalized (since gradients scale as $1/\tau$).
3.1 ADPO as a Unified Anchored Projection
We cast preference learning as an anchored projection on the probability manifold: given target $q$ and anchor $\pi_{\mathrm{ref}}$, minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ aligns the student under the softmax Fisher geometry. This unified view encompasses supervised fine-tuning, knowledge distillation, maximum-entropy RL, and DPO as special cases.
Master formulation.
@@ 3.1 ADPO as a Unified Anchored Projection
Given student $\pi_{\theta}$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau>0$, the anchored logits are
$u_{i}\;=\;\frac{s_{i}-s_{i}^{\mathrm{ref}}}{\tau},\qquad s_{i}=\log\pi_{\theta}(i\mid x),\ \ s_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(i\mid x).$ (3)
@@
Note (sequence models): In autoregressive sequence models, we use length-normalized log-probabilities (token-average NLL) for $s$ and $s^{\mathrm{ref}}$ to avoid length bias. Under this setting, each candidate’s score is the per-token average, and these averaged scores enter the group-level softmax over candidates. Since length normalization amounts to an equal rescaling of all scores within a group, it preserves the softmax affine-invariance: the Fisher geometry and anchoring properties remain intact when applied to this sequence-level distribution over candidates (not the token-level distribution). Thus the theoretical results—groupwise shift invariance (Lemma 2.1), implicit trust region (Lemma 3.3), and closed-form projection (Proposition 3.2)—carry through in this setting. The anchored model distribution over a group $S_{x}$ is
@@
$\tilde{p}_{\theta}(i\mid S_{x})\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}\;=\;\mathrm{softmax}_{i}\Big(\tfrac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau}\Big).$ (4)
Given a target preference distribution $q(\cdot\mid S_{x})$ (pairwise or listwise), ADPO minimizes
$\mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\left[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\right]\;=\;\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right]+\mathrm{const}.$ (5)
This objective performs a projection onto the anchored probability manifold; its second-order expansion around $q$ yields the softmax Fisher metric, inducing an implicit trust region centered at $\pi_{\mathrm{ref}}$ (formalized in Section 3.4).
Relation to standard DPO.
DPO [1] also introduces a reference model, but it enters only through log-odds differences between two candidates:
$\Delta_{\theta,\mathrm{ref}}=\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{i}}-\big(\log\pi_{\theta}-\log\pi_{\mathrm{ref}}\big)_{y_{j}}.$
In contrast, ADPO anchors each score $s_{i}-s_{i}^{\mathrm{ref}}$ individually before normalizing across the group to form $\tilde{p}_{\theta}$. When $|S_{x}|=2$ and $q$ is binary (or Bradley–Terry soft targets), Eq. (5) reduces to the logistic preference loss used in DPO. For general listwise $|S_{x}|>2$, ADPO retains distribution-level anchoring and inherits a Fisher-metric trust region around $\pi_{\mathrm{ref}}$.
@@
Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover:
@@
    Supervised fine-tuning (SFT). If $q(\cdot\mid S_{x})$ equals the data label distribution (one-hot or soft), $\pi_{\mathrm{ref}}$ is uniform (constant logits that cancel in softmax), and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss. For $\tau\neq 1$, $\tilde{p}_{\theta}$ is a temperature-scaled version of $\pi_{\theta}$.
@@
    Maximum-entropy RL / offline policy projection.
    • (Soft-optimal Boltzmann policy) The standard maximum-entropy RL objective yields the soft-optimal policy $\pi^{\star}(a|s)\propto\exp(Q(s,a)/\alpha)$ [19, 21, 20]. To match this target via ADPO’s anchored distribution $\tilde{p}_{\theta}$, we set $q(a|s)\propto\exp(Q(s,a)/\alpha)$ and use a uniform anchor $\pi_{\mathrm{ref}}$.
      Parametrization: ADPO constrains the anchored distribution $\tilde{p}(u)=\mathrm{softmax}((\log\pi_{\theta}-\log\pi_{\mathrm{ref}})/\tau)$ to equal $q$. Since there are equivalent parameterizations, we choose $\tau=1$ and $\pi_{\theta}(a|s)\propto\pi_{\mathrm{ref}}(a|s)\exp\!\big(Q(s,a)/\alpha\big)$ (with $\pi_{\mathrm{ref}}$ uniform, this reduces to $\pi_{\theta}\propto\exp(Q/\alpha)$). This directly recovers the standard soft-optimal policy $\pi^{\star}$ for use in sampling/evaluation, preserving the conventional Boltzmann temperature $1/\alpha$. Alternative parameterizations (e.g., $\tau=\alpha$ with $\pi_{\theta}\propto\exp(Q)$) yield the same $\tilde{p}$ but would shift the temperature into the anchoring scale, which we avoid for clarity.
    • (AWR/AWAC/MPO advantage-weighted projection) When $S_{x}$ is the full action space with overlapping support between $\pi_{\text{old}}$ and the target $q$, the anchor is $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, and the target is $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$ (shift-invariant to constant baselines), with $\tau=\beta$. The optimal anchored distribution is $\tilde{p}_{\theta}=q\propto\pi_{\text{old}}\exp(A/\beta)$, exactly matching the advantage-weighted target.
      Derivation of $\pi_{\theta}$ form: ADPO enforces $\tilde{p}(u)=q$ with $u=(\log\pi_{\theta}-\log\pi_{\text{old}})/\tau$ and $\tau=\beta$. Then
      $\tilde{p}(u)\ \propto\ \exp\!\Big(\frac{\log\pi_{\theta}-\log\pi_{\text{old}}}{\beta}\Big)\ \stackrel{{\scriptstyle!}}{{\propto}}\ \pi_{\text{old}}\,\exp\!\big(A/\beta\big),$
      which is equivalent to
      $\frac{\pi_{\theta}}{\pi_{\text{old}}}\ \propto\ \Big(\pi_{\text{old}}\,e^{A/\beta}\Big)^{\beta}\ \Rightarrow\ \pi_{\theta}\ \propto\ \pi_{\text{old}}^{\,1+\beta}\ e^{A}.$
      What is exactly recovered? ADPO matches the anchored distribution $\tilde{p}$ to the advantage-weighted target $q$; the induced policy shape $\pi_{\theta}\propto\pi_{\text{old}}^{\,1+\beta}\exp(A)$ is a consequence of the anchored projection (Prop. 3.2) and is not the classic MPO/AWAC M-step update (which yields $\pi_{\theta}\propto\pi_{\text{old}}\exp(A/\beta)$ or directly $\pi_{\theta}=q$ under sufficient capacity). Our equivalence is thus at the $\tilde{p}-$level rather than the $\pi_{\theta}$-level. Assumptions: (1) full action space $S_{x}$ is enumerable or covered by samples; (2) advantage weights $\exp(A/\beta)$ are normalizable; (3) for off-policy data $d^{\beta}$, we approximate by ignoring or clipping importance ratios (standard in AWR/AWAC [24, 25]), yielding weighted MLE $\max_{\theta}\mathbb{E}_{(s,a)\sim d^{\beta}}[\exp(A/\beta)\log\pi_{\theta}(a|s)]$. MPO’s E-step weights $w\propto\exp(Q/\eta)$ with dual variable $\eta$ (enforcing a KL constraint via Lagrangian duality) follow the same projection form [26], though MPO additionally solves for $\eta$ to satisfy a KL budget.

### v5

@@ 1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy-tailed label noise, DPO’s performance degrades by up to 71% relative to our method (0.472 vs. 0.809 WinMass). This sensitivity arises because the DPO objective only regularizes differences of log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms. Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3, Proposition 3.2). By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
We further show that this anchoring induces a principled geometric structure: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.5), yielding an implicit trust region centered at the target (not at the anchor). This offers a distribution-space analogue to TRPO [23] without explicit constraints. Because the geometry depends only on $q$, we can cleanly compare anchor update strategies. Under our experimental settings we find a task-dependent trade-off:
  • Dynamic anchors suit online exploration in noisy environments, improving performance by 5–11% in 3/4 scenarios on our benchmarks.
@@ 2 Related Work
Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the (unanchored) probability simplex. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates. Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint. We are not aware of prior work that formulates an anchored cross-entropy maintaining a fixed geometric center in probability space.
@@
$\max_{\pi}\ \mathbb{E}_{s\sim\mathcal{D}}\Big[\mathbb{E}_{a\sim\pi(\cdot|s)}r(s,a)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|s)\big)\Big],$
@@
$\pi^{*}(a|s)\;=\;\frac{1}{Z(s)}\exp\!\Big(\tfrac{1}{\alpha}\,r(s,a)\Big),$ (1)
@@
$r(s,a)\;=\;\alpha\,\log\pi^{*}(a|s)\;+\;\alpha\,\log Z(s),$ (2)
so any relative quantity (pairwise/listwise) cancels $Z(s)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs. (3)–(6), (20)–(27)).
@@
For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [18], Prop. 3.1 and App. D.2.
@@ 3 Anchored Direct Preference Optimization
$x$: context (prompt/state); $S_{x}=\{y_{1},\ldots,y_{|S_{x}|}\}$: candidate set (actions/outputs). $\pi_{\theta}(y\mid x)$: student policy; $\pi_{\mathrm{ref}}(y\mid x)$: anchor (fixed or dynamic). $s_{y}=\log\pi_{\theta}(y\mid x)$, $s^{\mathrm{ref}}_{y}=\log\pi_{\mathrm{ref}}(y\mid x)$: log-probabilities. $q(y\mid S_{x})$: target preference distribution (from teacher, rewards, or labels). $u_{y}=(s_{y}-s^{\mathrm{ref}}_{y})/\tau_{\mathrm{anc}}$: anchored logits. $\tilde{p}_{\theta}(\cdot\mid S_{x})=\mathrm{softmax}(u)$: anchored model distribution.
@@
$\tau_{\mathrm{anc}}$ Anchoring temperature Defines anchored logits: $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ (listwise)
@@
Remark (scale non-identifiability): In pairwise settings, temperatures $\beta$ and $\beta_{r}$ only appear through their ratio $\beta_{r}/\beta$ in the Bayes-optimal condition, so the statistical optimum is invariant to jointly rescaling them. We set $\beta=\beta_{r}=1.0$ by default. However, the optimization dynamics are affected: pairwise gradients scale as $\beta$ (since $\partial\ell/\partial\theta=\beta(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij})(\nabla s_{i}-\nabla s_{j})$), while listwise gradients scale as $1/\tau_{\mathrm{anc}}$ (because $\partial\mathcal{L}/\partial s_{i}=\tfrac{1}{\tau_{\mathrm{anc}}}(\tilde{p}-q)$). Thus smaller $\beta$ slows pairwise updates; smaller $\tau_{\mathrm{anc}}$ tightens the trust region via curvature $\propto 1/\tau_{\mathrm{anc}}^{2}$ but increases gradient magnitudes.
@@
Remark (temperature $\tau_{\mathrm{anc}}$ and regularization strength): In listwise settings with anchoring, $\tau_{\mathrm{anc}}$ controls the balance between the anchor and target via the closed-form solution $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ (Proposition 3.4). Intuitively:
@@
  • As $\tau_{\mathrm{anc}}\to 0$: $q^{\tau_{\mathrm{anc}}}\to\mathbf{1}$ (uniform), so $\pi_{\theta}\to\pi_{\mathrm{ref}}$ (stronger regularization toward the anchor).
@@
  • As $\tau_{\mathrm{anc}}\to\infty$: $q^{\tau_{\mathrm{anc}}}$ becomes increasingly peaked at $\operatorname*{arg\,max}q$, so $\pi_{\theta}$ tracks the target $q$ more aggressively.
@@
Note that, when the optimum is reachable, $\tilde{p}_{\theta}$ matches $q$ at optimality for any $\tau_{\mathrm{anc}}$; the temperature only controls how the underlying policy relates to the anchor via $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$. Moreover, locally around the optimum, a smaller $\tau_{\mathrm{anc}}$ induces a tighter implicit trust region: deviations in $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ are penalized more strongly since $\|\nabla_{s}\mathcal{L}\|\propto 1/\tau_{\mathrm{anc}}$ and the local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$.
3.1 Core Formulation: Anchored Distribution as the Protagonist
ADPO performs preference learning via anchored projection: we define a distribution $\tilde{p}_{\theta}$ in coordinates centered at a reference policy $\pi_{\mathrm{ref}}$, then minimize its KL divergence to a target $q$. All theory revolves around this anchored distribution $\tilde{p}_{\theta}$.
Definition of anchored distribution.
@@ 3.1 Core Formulation: Anchored Distribution as the Protagonist
Given student policy $\pi_{\theta}$, reference anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}>0$, we define:
  • Anchored logits:
    $u_{i}\;=\;\frac{\log\pi_{\theta}(i\mid x)-\log\pi_{\mathrm{ref}}(i\mid x)}{\tau_{\mathrm{anc}}}.$ (3)
  • Anchored distribution:
    $\tilde{p}_{\theta}(i\mid S_{x})\;=\;\mathrm{softmax}(u)_{i}\;=\;\frac{\exp(u_{i})}{\sum_{j\in S_{x}}\exp(u_{j})}.$ (4)
  • Objective: Given target distribution $q(\cdot\mid S_{x})$,
    $\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot\mid S_{x})\,\|\,\tilde{p}_{\theta}(\cdot\mid S_{x})\big)\Big].$ (5)
Remark 3.1 (Sequence-level scoring).
@@
For autoregressive models, we use length-normalized (per-token average) log-probabilities for $s$ and $s^{\mathrm{ref}}$ to avoid length bias. This preserves softmax affine-invariance, so all theoretical results (shift invariance, Fisher geometry, closed-form solutions) apply to the sequence-level distribution over candidates.
Groupwise shift invariance.
Since $\tilde{p}_{\theta}$ depends only on log-odds differences $u_{i}-u_{j}=(s_{i}-s_{i}^{\mathrm{ref}})-(s_{j}-s_{j}^{\mathrm{ref}})$ (via softmax), adding any constant $c$ to all $s_{i}$ and $s_{i}^{\mathrm{ref}}$ within a group leaves $\tilde{p}_{\theta}$ unchanged (Lemma 2.1). This cancels group-level biases—e.g., annotator-specific baselines—automatically.
Closed-form optimal solution.
When the anchored distribution $\tilde{p}_{\theta}$ can be matched exactly to the target $q$ (e.g., sufficient expressivity, full action coverage), the optimal policy takes the multiplicative fusion form (Proposition 3.4):
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (6)
@@
Proof sketch: $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ iff $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$, i.e., $\pi_{\theta}\propto\pi_{\mathrm{ref}}q^{\tau_{\mathrm{anc}}}$. ∎
Fisher geometry = implicit trust region.
The KL objective $\mathrm{KL}(q\|\tilde{p}_{\theta})$ admits a second-order expansion around the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$):
$\mathrm{KL}(q\|\tilde{p}_{\theta})\;\approx\;\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\operatorname{Var}_{q}\!\big[s-s^{\mathrm{ref}}-\tau_{\mathrm{anc}}\log q\big],$ (7)
where $\delta=u-u^{\star}$ is $q$-centered (Lemma 3.5). The matrix $\operatorname{Diag}(q)-qq^{\top}$ is the softmax Fisher metric, inducing a local quadratic trust region whose curvature depends only on $q$ (not on the anchor $\pi_{\mathrm{ref}}$). This provides a distribution-space analogue of TRPO’s trust region, but implicit—no explicit KL constraint needed.
DPO as a binary special case.
When $|S_{x}|=2$, the anchored distribution reduces to $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}})$ where $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$. With Bradley–Terry targets $q(i)=\sigma(\beta_{r}\Delta R)$ and uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$), Eq. (5) recovers the standard DPO loss [1]. For listwise $|S_{x}|>2$, ADPO generalizes to Plackett–Luce targets while retaining the Fisher-metric trust region.
Unification of learning paradigms.
Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning objectives as special cases:
@@
Let $\mathcal{L}_{\mathrm{ADPO}}$ be defined in Eq. (5). Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning objectives as special cases:
@@
    Supervised fine-tuning (SFT). When $q(\cdot\mid S_{x})$ is the data label distribution, $\pi_{\mathrm{ref}}$ is uniform, and $\tau_{\mathrm{anc}}=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, recovering standard cross-entropy loss.
@@
    Maximum-entropy RL. When $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and $\pi_{\mathrm{ref}}$ is uniform, setting $\tau_{\mathrm{anc}}=1$ recovers the soft-optimal Boltzmann policy $\pi_{\theta}\propto\exp(Q/\alpha)$ [19, 21].